This is a Solving by substution x+2y =8 and 2x+y =8
Answer by unlockmath(1688)

Hello,
Let's do this: x+2y =8 and 2x+y =8 Rewrite the first equation to look like:
x =8-2y and plug this into the second equation to get:
2(8-2y)+y=8 rewritten will be:
16-4y+y=8 Combine like terms and subtract 16 from both sides to get:
-3y=-8 Then:
y=8/3
Plug this answer into the original and we get:
x=8/3
Make sense?
There's a book I wrote on the basics
www.math-unlock.com